OpenStudy (anonymous):

Let f(x) = 3x+2 and g(x) = x2 Find f(g(x))

OpenStudy (zzr0ck3r):

so f(g(x)) = 3(g(x))+2 = 3(x^2)+2 = 3x^2+2

OpenStudy (anonymous):

thanks